OpenStudy (anonymous):

Identify the 42nd term of an arithmetic sequence where a1 = -12 and a27 = 66. 70 72 111 114

OpenStudy (anonymous):

we can get common difference d from a27 a27 = a + 26d 66 = -12 + 26*d 66+12 = 26d 26d = 78 d = 78/26 = 3 now a42 = a + 41d = -12 + 41*3 = -12 + 123 = 111 answer is 111